# Reassessing Active Learning Adoption in Contemporary NLP: A Community Survey

### v3

@@ Acknowledgments
We sincerely thank all the great anonymous participants of the survey for their time and invaluable answers, which enabled this research. Likewise we are grateful for everyone who supported us by distributing the survey, in particular Rabiraj Bandyopadhyay, David Berenstein, Ariel Gera, Alon Halfon, Philip H., Erik Körner, Gebriella Lapesa, Maximilian Maurer, Ines Montani, Paul Noirel, Julian Oestreich, Myrthe Reuver, Arij Riabi, Stefan Schweter, Artem Shelmanov, Eyal Shnarch, Manuel Tonneau and Akim Tsvigun. We are equally grateful to all anonymous supporters, also in cases in which we did not receive direct feedback on our requests for sharing. We would like to thank all tool authors and friends from the active learning community for their support. We sincerely appreciate the contributions of all who helped, even if not mentioned by name.
Moreover, we thank Ranjit Singh of GESIS Survey Methods Consulting for his expertise and feedback on early drafts of the survey, Harry Scells and Erik Körner who reviewed and tested the survey, Gerhard Heyer and Ingolf Römer for supporting the GDPR agreements, and Bildungsportal Sachsen for providing the infrastructure (LimeSurvey service). We thank Constantin Hammer, Luisa Köhler, and Jens Syckor for supporting us with the GDPR-compliance in their roles as data protection officers. We thank Ana Maria Lisboa dos Santos Cotovio for her support in the data collection for Figure 1.
This paper has been partially funded by the Federal Ministry of Education and Research of Germany and by Sächsische Staatsministerium für Wissenschaft, Kultur und Tourismus in the programme Center of Excellence for AI-research “Center for Scalable Data Analytics and Artificial Intelligence Dresden/Leipzig”, project identification number: ScaDS.AI. Further parts of this work have been conducted within the CORAL project which was funded by the Federal Ministry of Education and Research (BMBF) under the grant number 01IS24077A. Responsibility for the content of this publication lies with the authors.
@@
Appendix A Active Learning Publications (Figure 1)
We start from the full ACL anthology as BibTeX, including abstracts (January 31st, 2025). In a first step, only publications published in the context of AACL, ACL, COLING, EACL, EMNLP, NAACL, CL, and TACL between 2009 and 2024 were kept. This includes affiliated workshops. The resulting number of papers constitutes the total number of publications. Subsequently, we filtered publications that contain “active” in title or abstract. If “active” is directly followed by “learning”, the publication is kept without further checks, in case of other next words, we manually checked the publication’s relevance. The resulting number of papers constitutes the total number of AL publications.
@@ B.1 Survey Distribution and Advertisement
We chose a diverse distribution strategy, covering mailing lists, direct contacts, social media, and service providers. While some overlap between these channels may exist, we utilize all of them to ensure a wide outreach. Our research interest is on applications for text data, thus, we advertise the survey with a focus on the NLP community.
@@
The ACL mailing list covers all current members of the Association of Computational Linguistics – about 10,400 as of end of 2024. The worldwide recipients are researchers from academia and industry likewise. [^10] The call to the ACL mailing list was sent out only once, due to two reasons: first, a manual inspection period of about two weeks by the administrators before distribution limited us to requesting only two mailings during the survey period; second, our second request was deleted for unclear reasons. The ELRA corpora-list is managed by the European Language Resources Association. It serves as a platform for sharing information related to linguistic resources, therefore providing a way to reach researchers and practitioners involved in data annotation in their daily work. The exact number of recipients on the mailing list is not publicly available; however, it is widely assumed as reaching a substantial audience. The members of the ELRA SIGUL-list are professionals involved in the development of language resources and technologies for under-resourced languages. By utilizing this mailing list, we aim to gain a more diverse perspective on supervised learning and AL for lesser-studied languages, trying to mitigate potential biases in data collection driven by the uneven availability of language resources. The exact number of recipients on the mailing list is not publicly available. The Natural Language Processing Data Community is a google group [^11] https://groups.google.com/a/datacommunitydc.org/g/nlp for anyone interested in NLP including computational linguists, data scientists, and software engineers. It counts 341 members as of the end of 2024. The tada.cool initiative [^12] https://sites.google.com/view/polsci-ml-initiative/talks is a community of interdisciplinary researchers, with a focus on machine learning and NLP for social science applications. We reached 748 members via the slack channel. We also requested distribution via the linguistlist [^13] https://linguistlist.org/submit/qs/, but did not receive approval.
@@
We curated a list of researchers who co-authored one or more publications in major NLP venues over the past 15 years with AL as the central topic. To compile this list of experts, we started from the AL publications identified as described in Appendix A. We manually extracted the email addresses from all publications that mentioned AL directly in the title. The approach also includes papers from workshops to capture more researchers that have worked on practical application use cases. We contacted all 601 identified individuals by a personalized mail. In 139 cases, the extracted e-mail addresses were no longer functional. We added 9 existing personal contacts to complement the list of active researchers. These were likewise contacted per mail.
@@
To identify relevant tools, we supplemented a recent survey of annotation tools (Borisova et al., 2024) with additional tools we encountered through years of research in AL. We contacted: ActiveAnno Wiechmann et al. (2021), AL Toolbox Tsvigun et al. (2022b), Argilla [^14] https://github.com/argilla-io/argilla, AWS Sagemaker Ground Truth and Comprehend, BioQRator Kwon et al. (2013), CleanLab [^15] https://github.com/cleanlab/cleanlab, INCEpTION Klie et al. (2018), Labelbox [^16] https://labelbox.com/, Label Sleuth Shnarch et al. (2022a), Label Studio [^17] https://labelstud.io/, MITRE Annotation Toolkit (MAT) [^18] https://sourceforge.net/projects/mat-annotation/, and POTATO Pei et al. (2022). We also contacted several companies and services via their official communication channels, asking them to forward the survey to their users. As an incentive, we highlighted the potential insights that could be gained into the actual needs of users in supervised learning and AL. Although this approach yielded limited feedback only, it nevertheless contributed to the overall outreach. Positive responses were received from AL Toolbox, which forwarded the call to some of their annotators; Argilla, which permitted the call to be shared in the Hugging Face Discord channel; Explosion, which shared the call on LinkedIn; Labelbox, which forwarded the call internally; and Label Sleuth, which forwarded the call to their users.
@@ B.4 Data Protection
In consultation with data protection officers, we implemented measures to ensure compliance with the EU General Data Protection Regulation (GDPR). Following the GDPR principle of data minimization, we collected only data necessary for our research objectives. This excluded, e.g., information on how long users took to complete the survey.
The survey was conducted anonymously, with the option to provide an email address to receive the survey evaluation afterward. Free-form text fields constitute a gray area. This is because participants could potentially enter personally identifiable information, particularly when combined with the demographic data collected in question group VII. To meet legal obligations, we informed participants about how any potential personal data would be processed and concluded a joint-controller agreement among the institutions involved.
@@ B.5 Anonymization for Dataset Release
For public release, text answers were manually processed as follows. We started from the raw dataset, excluding the separately stored email addresses.
For each field, we replaced all mentions that could reveal the respondents’ identities with placeholders, including URLs, links, email addresses, single words, sentences, and paragraphs. We also applied this to language mentions, as they could contribute to de-anonymization when combined with other fields (e.g., in case of languages that few people in the NLP community are researching). In this context, we took a conservative approach to V.2 (languages of the annotation project with AL) by removing all answers to preserve anonymity.
For question items containing highly specific information that may allow to draw conclusions about the respective participant, we replaced the entries with broader categories: IV.1 (year): 2020–2024, 2015–2019, 2010–2014, 2005–2009; V.3 (hours): up to 20, 21–50, 51–100, 101–500, more than 500; V.4 (instance count): up to 100, 101–500, 501–1000, 1001–10,000, more than 10,000, N/A; V.4 (instance type): documents, sentences, tokens, other, N/A; VII.5 (countries): Africa, Asia, Europe, North America, Oceania, South America.
As a final step, we separated responses and demographic information and shuffled rows randomly to prevent re-merging of the two files.
@@ Appendix C Full Survey
The survey is initiated by the researchers [anonymized for review]. If you have any questions, please contact us via [anonymized for review].
@@
[not disclosed to ensure anonymity in the review]
@@ Appendix D Results
Table LABEL:tab:responses-full provides a full overview of survey participants’ responses to the predefined options for all 52 questions. It is beyond the scope of this paper to cover all analyses enabled by the collected data, but the public release of the dataset will allow researchers to conduct in-depth analyses of aspects they are particularly interested in.
@@ D.1 Conservative Estimates (No Authors)
Our targeted recruitment of authors may have led to potential overestimation of AL in question groups I and II (the non-AL parts of the survey). To transparently account for potential author bias (response bias), our data includes information on recruitment methods (see VII.7). By excluding the 33 participants recruited via personalized mails, we can provide a conservative baseline. Explicitly excluding authors of AL provides the following picture:
@@
Dissemination of active learning AL was used by 36 respondents (40%; vs. 53% in the full sample). When asking why participants had not chosen AL, 25% (vs. 25% in the full sample) had never heard of it, and 51% (vs. 49% in the full sample) cited insufficient expertise. Methodological concerns cover the expected implementation overhead (40%; vs. 37% in the full sample), unsuitable annotation tools (34%; vs. 32% in the full sample), the difficulty of estimating effectiveness upfront (21%; vs. 12% in the full sample), and sampling bias (15%; vs. 18% in the full sample). 9% (vs. 9% in the full sample) found it unsuitable due to specific project requirements, and 26% (vs. 14% in the full sample) doubted AL’s effectiveness at all.
@@
Alternative methods The current users of AL also report obstacles in data annotation with AL. These are: finding annotators (36%; vs. 35% in the full sample), AL with multiple annotators (53%; vs. 56% in the full sample), unclear annotation instructions (31%; vs. 29% in the full sample), hard cases for annotation (53%; vs. 59% in the full sample), lack of GPU/TPU resources (22%; vs. 25% in the full sample), choice of model / query strategy (44%; vs. 49% in the full sample), choice of stopping criterion (36%; vs. 35% in the full sample), and other (6%; vs. 6% in the full sample). The response distributions closely resemble each other.
@@
Both groups show similar preferences for LLMs, with about two-thirds in each group using LLMs, primarily smaller models such as BERT. Industry/governmental participants rely more on uncertainty sampling (50% vs. 25%), and their waiting times vary more—possibly due to a larger data pool.
@@ D.3 Selected Additional Analysis
Waiting times
Figure 8 illustrates the waiting times between annotation cycles in practice (reported for the 33 recent projects), and how this relates to the general perception of what is acceptable (reported by all AL users, lab and applied, in III.4). We observe a broad range reported across projects, from under one minute to up to one week. At the same time, participants considered even longer waiting times to be acceptable in practice, suggesting that significant sampling times are neither unexpected nor problematic for them.
Figure 8: Maximum acceptable as opposed to last encountered waiting times (if applicable).

### v4

@@ Acknowledgments
We thank all the great anonymous participants of the survey for their time and invaluable answers, which enabled this research. Likewise, we are grateful for everyone who supported us by distributing the survey, in particular Rabiraj Bandyopadhyay, David Berenstein, Ariel Gera, Alon Halfon, Philip H., Erik Körner, Gebriella Lapesa, Maximilian Maurer, Ines Montani, Paul Noirel, Julian Oestreich, Myrthe Reuver, Arij Riabi, Stefan Schweter, Artem Shelmanov, Eyal Shnarch, Manuel Tonneau and Akim Tsvigun. We are equally grateful to all anonymous supporters. Thanks to all tool authors and friends from the AL community for their support. We sincerely appreciate the contributions of all who helped, even if not mentioned by name.
We thank Ranjit Singh of GESIS Survey Methods Consulting for his expertise and feedback on early drafts of the survey, Harry Scells and Erik Körner who reviewed and tested the survey, Gerhard Heyer and Ingolf Römer for supporting the GDPR agreements, and Bildungsportal Sachsen for providing the LimeSurvey infrastructure, Constantin Hammer, Luisa Köhler, and Jens Syckor for supporting us with the GDPR-compliance in their roles as data protection officers, and Ana Lisboa Cotovio for support in the data collection for Figure 1.
Parts of this paper have been financially supported by the German Federal Ministry of Research, Technology and Space (BMFTR) and by Sächsische Staatsministerium für Wissenschaft, Kultur und Tourismus in the programme Center of Excellence for AI-research “Center for Scalable Data Analytics and Artificial Intelligence Dresden/Leipzig”, project identification number: ScaDS.AI. Further parts of this work have been conducted within the CORAL project which was funded by the German Federal Ministry of Research, Technology, and Space (BMFTR) under the grant number 16IS24077A. Responsibility for the content of this publication lies with the authors.
@@
Appendix A Active Learning Publications
To obtain the data for Figure 1, we start from the full ACL anthology in BibTeX format, including abstracts (January 14th, 2026). In a first step, only publications published in the context of AACL, ACL, COLING, EACL, EMNLP, NAACL, CL, and TACL between 2009 and 2025 were kept. This includes affiliated workshops. The resulting number of papers constitutes the total number of publications. Subsequently, we filtered publications that contain “active” in title or abstract. If “active” is directly followed by “learning”, the publication is kept without further checks, in case of other next words, we manually checked the publication’s relevance. [^10] This way, we account for wording variations such as “active fine-tuning” or “active sampling.” The resulting number of papers constitutes the total number of AL publications.
The numbers for 2025 must be interpreted with caution, as the ACL Anthology had not yet been fully updated with all entries at the time of analysis (e.g., the proceedings of AACL, some workshops).
@@ B.1 Survey Distribution and Advertisement
We chose a diverse distribution strategy, covering mailing lists, direct contacts, social media, and service providers. While some overlap between these channels may exist, we utilize all of them to ensure a wide outreach.
@@
The ACL mailing list covers all current members of the Association of Computational Linguistics – about 10,400 as of end of 2024. The worldwide recipients are researchers from academia and industry likewise. [^11] The call to the ACL mailing list was sent out only once, due to two reasons: first, a manual inspection period of about two weeks by the administrators before distribution limited us to requesting only two mailings during the survey period; second, our second request was deleted for unclear reasons. The ELRA corpora-list is managed by the European Language Resources Association. It serves as a platform for sharing information related to linguistic resources, therefore providing a way to reach researchers and practitioners involved in data annotation in their daily work. The exact number of recipients on the mailing list is not public; however, it is widely assumed as reaching a substantial audience. The members of ELRA SIGUL-list are professionals involved in the development of language resources and technologies for under-resourced languages. By utilizing this mailing list, we aim to gain a more diverse perspective on supervised learning and AL for lesser-studied languages, trying to mitigate potential biases in data collection driven by the uneven availability of language resources. The exact number of recipients on the mailing list is not public. Natural Language Processing Data Community is a google group [^12] https://groups.google.com/a/datacommunitydc.org/g/nlp for anyone interested in NLP including computational linguists, data scientists, and software engineers. It counts 341 members as of the end of 2024. The tada.cool initiative [^13] https://sites.google.com/view/polsci-ml-initiative/talks is a community of interdisciplinary researchers, with a focus on machine learning and NLP for social science applications. We reached 748 members via the slack channel.
@@
We curated a list of researchers who co-authored one or more publications on AL in major NLP venues over the past 15 years. To compile this list of experts, we started from the AL publications identified as described in Appendix A. We manually extracted the email addresses from publications that mentioned AL directly in the title. The approach also includes papers from workshops to capture more practical application experiences. We contacted all 601 identified individuals by a personalized mail. In 139 cases, the extracted e-mail addresses were no longer functional. We added 9 further personal contacts to complement the list.
@@
To identify relevant tools, we supplemented a recent survey of annotation tools (Borisova et al., 2024) with additional tools we encountered through years of research in AL. We contacted: ActiveAnno Wiechmann et al. (2021), AL Toolbox Tsvigun et al. (2022b), Argilla [^14] https://github.com/argilla-io/argilla, AWS Sagemaker Ground Truth and Comprehend, BioQRator Kwon et al. (2013), CleanLab [^15] https://github.com/cleanlab/cleanlab, INCEpTION Klie et al. (2018), Labelbox [^16] https://labelbox.com/, Label Sleuth Shnarch et al. (2022a), Label Studio [^17] https://labelstud.io/, MITRE Annotation Toolkit (MAT) [^18] https://sourceforge.net/projects/mat-annotation/, and POTATO Pei et al. (2022). We also contacted several companies and services via their official communication channels, asking them to forward the survey to their users. As an incentive, potential insights that could be gained into the actual needs of users in supervised learning and AL were emphasized. Although this approach yielded limited feedback, it nevertheless contributed to the overall outreach. Positive responses were received from AL Toolbox, which forwarded the call to some of their annotators; Argilla, which permitted the call to be shared in the Hugging Face Discord channel; Explosion, which shared the call on LinkedIn; Labelbox, which forwarded the call internally; and Label Sleuth, which forwarded the call to their users.
@@ B.4 Data Protection
In consultation with data protection officers, we implemented measures to ensure compliance with the EU General Data Protection Regulation. Following it’s principle of data minimization, we collected only data necessary for our research objectives. The survey was conducted anonymously, with the option to provide an email address to receive the survey evaluation afterward. Free-form text fields are a gray area, as participants may potentially enter personally identifiable information, particularly when combined with demographic data collected. To meet legal obligations, we informed participants about how any personal data would be processed, and concluded a joint-controller agreement among the institutions involved.
@@ B.5 Anonymization for Dataset Release
For public release, text answers were manually processed as follows. We started from the raw dataset, excluding the separately stored email addresses. For each field, we replaced all mentions that could reveal the respondents’ identities with placeholders, including URLs, links, email addresses, single words, sentences, and paragraphs. We also applied this to language mentions, as they could contribute to de-anonymization when combined with other fields (e.g., in case of languages that few people in the NLP community are researching). In this context, we took a conservative approach to V.2 and V.7 by removing all answers to preserve anonymity. For question items containing highly specific information that may allow to draw conclusions about the respective participant, we replaced the entries with broader categories: IV.1 (year): 2020–2024, 2015–2019, 2010–2014, 2005–2009; V.3 (hours): up to 20, 21–50, 51–100, 101–500, more than 500; V.4 (instance count): up to 100, 101–500, 501–1000, 1001–10,000, more than 10,000, N/A; V.4 (instance type): documents, sentences, tokens, other, N/A; VII.5 (countries): Africa, Asia, Europe, North America, Oceania, South America.
Responses and demographics are released separately and randomly shuffled to prevent re-merging.
@@ Appendix C Full Survey
The survey is initiated by the researchers Christopher Schröder (Institut für Angewandte Informatik e. V.), Julia Romberg (GESIS - Leibniz Institute for the Social Sciences), and Julius Gonsior (TU Dresden). If you have any questions, please contact us via activelearningsurvey2024@gmail.com.
@@
[link to declaration of consent]
@@ Appendix D Results
Table LABEL:tab:responses-full provides a full overview of survey participants’ responses to the predefined options for all 52 questions. It is beyond the scope of this paper to cover all analyses enabled by the collected data, but the public release will allow researchers to conduct in-depth analyses of further aspects.
@@ D.1 Conservative Estimates (No Authors)
Our targeted recruitment of authors may have led to potential overestimation of AL in question groups I and II (the non-AL parts of the survey). To transparently account for potential author bias (response bias), our data includes information on recruitment methods (see VII.7). By excluding the 33 participants recruited via personalized mails, we can provide a conservative baseline. Excluding authors of AL publications provides the following picture:
@@
Adoption of active learning: AL was used by 36 respondents (40%; vs. 53% in the full sample). When asking why participants had not chosen AL, 25% (vs. 25% in the full sample) had never heard of it, and 51% (vs. 49% in the full sample) cited insufficient expertise. Methodological concerns cover the expected implementation overhead (40%; vs. 37% in the full sample), unsuitable annotation tools (34%; vs. 32% in the full sample), the difficulty of estimating effectiveness upfront (21%; vs. 12% in the full sample), and sampling bias (15%; vs. 18% in the full sample). 9% (vs. 9% in the full sample) found it unsuitable due to specific project requirements, and 26% (vs. 14% in the full sample) doubted AL’s effectiveness at all.
@@
Alternative methods: The current users of AL also report obstacles in data annotation with AL. These are: finding annotators (36%; vs. 35% in the full sample), AL with multiple annotators (53%; vs. 56% in the full sample), unclear annotation instructions (31%; vs. 29% in the full sample), hard cases for annotation (53%; vs. 59% in the full sample), lack of GPU/TPU resources (22%; vs. 25% in the full sample), choice of model / query strategy (44%; vs. 49% in the full sample), choice of stopping criterion (36%; vs. 35% in the full sample), and other (6%; vs. 6% in the full sample). The response distributions closely resemble each other.
@@
Both groups show similar preferences for LLMs, with about two-thirds in each group using them, primarily the smaller models. Industry/governmental participants rely more on uncertainty sampling (50% vs. 25%), and their waiting times vary more—possibly due to a larger data pool.